# Priors in Time: Missing Inductive Biases for Language Model Interpretability

### v2

@@ G.1 Priors on the Sparse Code for various SAEs
Proposition G.1 (Independence priors over time).
@@
G.2 Sparsity over time and implications
@@
Corollary G.1.1 (Assumptions of time-invariant sparsity).
@@
Proposition G.2 (Restrictive Sparsity Budget Leads to Support Switching in SAEs).
@@
Appendix H Further Theoretical Results: SAE-wise Characterization of Priors over concepts and time
@@
Proposition H.1 (SAE Priors on Sparse Code).
@@ H.0.2 TopK SAE
As defined in Theorem H.1 (and restated here for convenience), let $S_{i}=\mathrm{supp}(\bm{z}_{i})=\{k:z_{i}^{k}>0\},n_{i}=|S_{i}|=\|\bm{z}_{i}\|_{0}$ denote the active indices and their number (sparsity) respectively.
@@ H.0.3 BatchTopK SAE
Following similar analysis as for TopK SAE (App. H.0.2), we can derive an equivalent prior over $\bm{z}$ for BatchTopK SAE:

### v3

@@ H.1 Priors on the Sparse Code for various SAEs
Proposition H.1 (Independence priors over time).
@@
H.2 Sparsity over time and implications
@@
Corollary H.1.1 (Assumptions of time-invariant sparsity).
@@
Proposition H.2 (Restrictive Sparsity Budget Leads to Support Switching in SAEs).
@@
Appendix I Further Theoretical Results: SAE-wise Characterization of Priors over concepts and time
@@
Proposition I.1 (SAE Priors on Sparse Code).
@@ I.0.2 TopK SAE
As defined in Theorem I.1 (and restated here for convenience), let $S_{i}=\mathrm{supp}(\bm{z}_{i})=\{k:z_{i}^{k}>0\},n_{i}=|S_{i}|=\|\bm{z}_{i}\|_{0}$ denote the active indices and their number (sparsity) respectively.
@@ I.0.3 BatchTopK SAE
Following similar analysis as for TopK SAE (App. I.0.2), we can derive an equivalent prior over $\bm{z}$ for BatchTopK SAE: